# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ Abstract
This study investigates a decentralized linear-quadratic optimal control problem, and several approximate separable constrained optimization problems are formulated for the first time based on the selection of sparsity promoting functions. First, for the optimization problem with weighted $\ell_{1}$ sparsity promoting function, a two-timescale algorithm is adopted that is based on the BSUM (Block Successive Upper-bound Minimization) framework and a differential equation solver. Second, a piecewise quadratic sparsity promoting function is introduced, and the induced optimization problem demonstrates an accelerated convergence rate by performing the same two-timescale algorithm. Finally, the optimization problem with $\ell_{0}$ sparsity promoting function is considered that is nonconvex and discontinuous, and can be approximated by successive coordinatewise convex optimization problems.
Keywords: decentralized linear-quadratic optimal control, separable optimization problem, two-timescale algorithm, $\ell_{0}$ sparsity promoting function.
@@ 1 Introduction
During the last few decades, the concept of decentralized control has gained significant attention due to its appealing characteristics, such as the potential to lower communication costs and parallelizable processing. In particular, decentralized linear-quadratic (LQ) optimal control is also a hot topic, which has been extensively investigated [25, 26, 7, 29]. Nevertheless, decentralized LQ problems present a significant challenge, as obtaining a closed solution is exceedingly difficult, with the exception of highly specific cases [46]. Optimization, an “end-to-end” method allowing us to optimize the performance criterion directly, has significantly contributed to the centralized LQ problems [11, 35, 16, 9, 8, 38, 10, 53] and decentralized LQ problems [52, 13, 28, 49], respectively. However, the question of how to design a controller with a general sparsity structure has yet to be resolved. Here, the term “sparsity structure” in this context pertains to the presence of zero elements within the stabilizing linear feedback gain matrix. Motivated by recent achievements in optimization-based LQ controls, the objective of this study is to establish a convex optimization framework for decentralized control incorporating a sparsity promoting penalty, which allows us to obtain an feedback gain with arbitrary sparsity structure.
@@ 1.1 Related Works
On the other hands, a canonical decentralized LQ optimal control can be cast as a centralized problem where the stabilizing feedback gain matrix is restricted to lie in a particular subspace $\mathcal{K}$. If $\mathcal{K}$ denotes the block-diagonal matrix space, classic convex parametrization methods can be utilized to obtain an approximate convex problem [9], and ADMM (Alternating Direction Method of Multipliers) based method is shown to converge to the global optimizer of the abovementioned approximate convex problem [52, 49, 24]. If $\mathcal{K}$ possesses general sparsity structure, the SDP (Semidefinite Programming) relaxation perspective is of concern. If the SDP relaxation has a rank-$1$ solution, a globally linear optimal distributed controller can be recovered from the solution of SDP [13]. Additionally, zero-order policy optimization algorithm is utilized to optimize the decentralized LQ cost directly, and it can converge to a stationary point with a PAC (Probably Approximately Correct) guarantee [28].
2) Sparse optimization. Sparse optimization, a task that is intrinsically NP-hard, is a fundamental challenge in the fields of optimization, machine learning, image processing, and control theory. One popular method is the relaxed approximation method, for example, convex relaxed method [4, 51, 44], nonconvex relaxed method [48, 15] and DC (difference of convex functions) relaxed method [21, 27]. Due to an enhanced comprehension of the variational characteristics associated with the $\ell_{0}$ norm, numerous studies have addressed sparse optimization by directly optimizing the $\ell_{0}$ norm [30, 2]. Additionally, successive coordinatewise convex problem can approximate $\ell_{0}$ norm penalty problem [42].
@@ 1.2 Contributions
  1. 1.
    Via the well-established concept of classic parameterization, we formulate a constrained optimization problem with a sparsity promoting penalty for the first time. Under our framework, the optimal linear feedback gain of a general decentralized LQ problem can be represented by the global minimizers of this issue in a certain sense.
  2. 2.
    Convex optimization framework introduced in [31] is feasible for the formulated optimization problem, provided that the sparsity promoting function is weighted $\ell_{1}$ norm. However, the framework of [31] cannot be utilized directly due to the presence of two distinct subproblems. The closed form solution of one subproblem is easily derived, while the remaining subproblems is complicated, necessitating the development of an iterative algorithm based on BSUM framework (see Algorithm 4). Observe that the process of BSUM entails the sequential resolution of various optimization issues, which have predefined structures yet, necessitating the need for individual design. Via the special and ingenious construction, each issue in BSUM possesses a closed form solution and BSUM exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate indeed.
  3. 3.
    Next, it is established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function. To our best knowledge, this phenomenon has not been reported in the filed of sparse optimization.
  4. 4.
    The variational properties of $\ell_{0}$-penalty is firstly studied for decentralized LQ problem. Motivated by the convex relaxation framework of [42], we show that the $\ell_{1}$-penalty induced problems can be seen as approximations of $\ell_{0}$-penalty induced problem in the context of decentralized LQ optimal control. Noting that [42] studies an unconstrained smooth optimization problem, the considered problem of this paper is constrained and nonsmooth. Hence, the conclusion and techniques of proof presented in our analysis do not exhibit parallel generalizations of those of [42].
The interconnection between the proposed optimization problems is illustrated in the diagram below.
[…]
@@
[…]
For any proper closed convex function $f$ on $\mathcal{X}$, we say $f\in\mathcal{S}_{\mu}^{0}(\mathcal{X})$ with $\mu\geq 0$ if
$f(x_{1})-f(x_{2})-\langle p,x_{1}-x_{2}\rangle\geq\frac{\mu}{2}\|x_{1}-x_{2}\|^{2},\leavevmode\nobreak\ \forall(x_{1},x_{2})\in\mathcal{X}\times\mathcal{X},$
where $p\in\partial f(x_{2})$ with $\partial f(x_{2})$ the subdifferential of $f$ at $x_{2}\in\mathcal{X}$. We write $f\in\mathcal{S}_{\mu,L}^{1,1}(\mathcal{X})$ if $f\in\mathcal{S}_{\mu}^{0}(\mathcal{X})$ has $L$-Lipschitz continuous gradient:
$f(x_{1})-f(x_{2})-\langle\nabla f(x_{2}),x_{1}-x_{2}\rangle\leq\frac{L}{2}\|x_{1}-x_{2}\|^{2},\leavevmode\nobreak\ \forall(x_{1},x_{2})\in\mathcal{X}\times\mathcal{X}.$
Given
$v=\begin{pmatrix}v_{11}&\cdots&v_{1n}\\ \vdots&\ddots&\vdots\\ v_{m1}&\cdots&v_{mn}\end{pmatrix}\in\mathbb{R}^{m\times n},$
the function ${\rm vec}(v):\mathbb{R}^{m\times n}\longmapsto\mathbb{R}^{mn}$ is introduced as
${\rm vec}(v)=(v_{11},\dots,v_{1n},\dots,v_{m1},\dots,v_{mn})^{\top}.$
For any $\alpha\in\mathbb{R}$, the $\alpha$-level set of a function $f:\mathbb{R}^{n}\to[-\infty,\infty]$ is the set
${\rm lev}_{\leq\alpha}f=\{x\in\mathbb{R}^{n}\colon f(x)\leq\alpha\}.$
We introduce the notation $M\lesssim N$, meaning $M\leq CN$ with some generic bounded constant $C>0$ that is independent of $A,B,\mu_{f},\mu_{g},\gamma_{0},\beta_{0}$ (see Section 5) but can be different in each occurrence.
@@ 2 Formulation
We introduce the following notation
$F=\begin{bmatrix}A&B_{2}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p},\leavevmode\nobreak\ G=\begin{bmatrix}0\\ I_{m}\end{bmatrix}\in\mathbb{R}^{p\times m},$
$Q=\begin{bmatrix}B_{1}B_{1}^{\top}&0\\ 0&0\end{bmatrix}\in\mathbb{S}^{p},\leavevmode\nobreak\ R=\begin{bmatrix}C^{\top}C&0\\ 0&D^{\top}D\end{bmatrix}\in\mathbb{S}^{p}.$
@@
$W=\begin{bmatrix}W_{1}&W_{2}\\ W_{2}^{\top}&W_{3}\end{bmatrix}\in\mathbb{S}^{p},$
where $W_{1}\in\mathbb{S}_{++}^{n},W_{2}\in\mathbb{R}^{n\times m},W_{3}\in\mathbb{S}^{m}$. Regard $\Theta_{i}(W)=F_{i}W+WF_{i}^{\top}+Q$ as a block matrix, i.e.,
@@
The following theorem introduces a subset of stabilizing controller gains, and find a upper bound of performance criterion (2).
@@
and let
$\mathscr{K}=\{K=W_{2}^{\top}W_{1}^{-1}\colon W\in\mathscr{C}\}.$
Then,
@@
    $K\in\mathscr{K}$ stabilizes the closed-loop system in the presence model uncertainties
@@
    $\langle R,W\rangle\geq\|H_{i}(s)\|_{\mathcal{H}_{2}}^{2},$
@@
Proof.
The proof is familiar with that of Theorem 1 of [32], but for the sake of completeness, we will include short proof here. For statement 1, $\Theta_{i,1}\preceq 0$ implies that
$(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})W_{1}+W_{1}(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})^{\top}+B_{1}B_{1}^{\top}\preceq 0.$ (3)
@@
$\langle R,W\rangle={\rm Tr}(CW_{1}C^{\top}+DW_{3}D^{\top})\geq{\rm Tr}(CW_{c,i}C^{\top}+DKW_{c,i}K^{\top}D^{\top})={\rm Tr}((C-DK)W_{c,i}(C-DK)^{\top}).$
The proof is complete. ∎
@@
The above theorem introduce a subset $\mathscr{K}$ of stabilizing feedback gain matrices and the upper bounded of performance criterion (2) in the presence of parameter uncertainties. Let us denote all the stabilizing feedback gain matrices as $\mathcal{S}$, and the we intend to solve the following optimization problem originally
$\min\limits_{K\in\mathcal{S}}J(K)={\rm Tr}\left((C-DK)W_{c}(C-DK)^{\top}\right),$ (4)
which is a constrained nonconvex problem. However, by the above theorem, we can solve the following optimization problem to approximate a “good” solution of (4)
$\min\limits_{W\in\mathscr{C}}\langle R,W\rangle,$ (5)
which is a convex problem since $\mathscr{C}$ is a convex set and objective function is linear. Thus, it seems that we obtain a “good” solution of a nonconvex problem by solving a convex optimization problem, and we have achieved substantial advancements. However, it is known that $J(K)$ is $L$-smooth and satisfies PL condition, so the gradient-based method can be utilized to obtain global minimum of (4) [10]. But if there are some sparsity constraints on $K$, the transformation of optimization problem from (4) to (5) is revolutionary, since the current state of scholarly knowledge remains uncertain regarding whether the gradient-based method could converge to the global minimum for the PL function with equality constraints.
In this study, we focus on a broad decentralized control problem, assuming that the feedback gain exhibits sparsity rather than a block diagonal structure [32]. The fundamental concept underlying our proposal is that the sparse stabilizing feedback gain, regardless of its structure, can be achieved by the design of a diagonal matrix $W_{1}$ and a sparse matrix $W_{2}$ with some sparse structure. Then, feedback gain $K=W_{2}^{\top}W_{1}^{-1}$ can preserve the sparsity structure of $W_{2}$ and stabilizes the closed-loop system.
@@
In this paper, we will use the weighted-$\ell_{1}$ norm, a piecewise quadratic function, and the $\ell_{0}$ norm as sparsity promoting functions, which correspond to distinct optimization problems. First, for the weighted-$\ell_{1}$ norm of a matrix $F$
$g(F)=\sum_{i,j}w_{ij}|F_{ij}|,$ (6)
introduce the following optimization problem
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$ (7)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $W\in\mathbb{S}^{p}_{+},$
$-V_{2}(F_{i}W+WF_{i}^{\top}+Q)V_{2}^{\top}\in\mathbb{S}_{+}^{n},\forall i=1,\dots,M,$
$-V_{j1}WV_{j2}=0,\forall j=1,\dots,\frac{n(n-1)}{2},$
$V_{1}WV_{2}^{\top}-P=0$
with $\gamma>0$, $V_{1}=[0,I_{m}],V_{2}=[I_{n},0]$. Denote
@@
Then, optimization problem (7) can be equivalently expressed in a compact form
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$ (8)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{G}(W)\in\mathcal{K},$
@@
and convex cone $\mathcal{K}$ is defined as
$\mathcal{K}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times 0\times\cdots\times 0.$
Since the positive semi-definite cone is self-dual, it is straightforward to express the dual cone $\mathcal{K}^{*}$ as
$\mathcal{K}^{*}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times\mathbb{R}\times\cdots\times\mathbb{R}.$

### v3

@@ Abstract
A $\mathcal{H}_{2}$-guaranteed decentralized linear-quadratic optimal control with convex parameterization and convex-bounded uncertainty is studied in this paper, where several sparsity promoting functions are added, respectively, into the $\mathcal{H}_{2}$ cost to penalize the number of communication links among decentralized controllers. Then, the sparse feedback gain is investigated to minimize the modified $\mathcal{H}_{2}$ cost together with the stability guarantee, and the corresponding main results are of three parts. First, the weighted-$\ell_{1}$ sparsity promoting function is of concern, and a two-timescale algorithm is developed based on the BSUM (Block Successive Upper-bound Minimization) framework and a primal-dual splitting approach. Second, the optimization problem induced by piecewise quadratic sparsity penalty is investigated, which exhibits an accelerated convergence rate. Third, the nonconvex sparse optimization problem with $\ell_{0}$-penalty is studied, which can be approximated by successive coordinatewise convex optimization problems.
{IEEEkeywords}
decentralized optimal control, linear-quadratic system, two-timescale algorithm, $\ell_{0}$-penalty
@@ 1 Introduction
During the last few decades, the concept of decentralized control has gained significant attention due to its appealing characteristics, such as the potential to lower communication costs and parallelizable processing. In particular, decentralized linear-quadratic (LQ) optimal control has been also a topic of general interest, which has been extensively investigated [3, 4]. Nevertheless, decentralized LQ problems present a significant challenge, as obtaining a closed-form solution is exceedingly difficult, with the exception of highly specific cases [5]. Optimization, an “end-to-end” method allowing us to optimize the performance criterion directly, has significantly contributed to the centralized LQ problems [6, 7, 8, 9, 10, 11, 12, 13] and decentralized LQ problems [14, 15, 16, 17], respectively. However, the question of how to design a controller with a general sparsity structure has yet to be resolved. Here, the term “sparsity structure” in this context pertains to the presence of zero elements within the stabilizing linear feedback gain matrix. In this paper, the $\mathcal{H}_{2}$-guaranteed cost decentralized control problem under convex parameterization and convex-bounded uncertainty is studied. Motivated by recent achievements in optimization-based LQ controls, the objective of this study is to establish a convex optimization framework for decentralized control incorporating a sparsity promoting penalty, allowing us to obtain a sparse feedback gain to lower the communication costs among controllers.
@@ 1.1 Related Works
On the other hands, a canonical decentralized LQ optimal control can be cast as a centralized problem where the stabilizing feedback gain matrix is restricted to lie in a particular subspace $\mathcal{K}$. If $\mathcal{K}$ denotes the block-diagonal matrix space, classic convex parametrization methods can be utilized to obtain an approximate convex problem [9], and ADMM (Alternating Direction Method of Multipliers) based method is shown to converge to the global optimizer of the abovementioned approximate convex problem [14, 17, 19]. If $\mathcal{K}$ possesses general sparsity structure, the SDP (Semidefinite Programming) relaxation perspective is of concern. If the SDP relaxation has a rank-$1$ solution, a globally linear optimal distributed controller can be recovered from the solution of SDP [15]. Moreover, an exact and inexact convex reformulation of sparse linear controller synthesis is introduced based on Youla parameterization and quadratically invariant assumption [41, 42], and infinite-dimensional convex optimization problems are proposed, which can be solved approximately by Laurent expansion. Additionally, zero-order policy optimization algorithm is utilized to optimize the decentralized LQ cost directly, and it can converge to a stationary point with a PAC (Probably Approximately Correct) guarantee [16].
2) Sparse optimization. Sparse optimization, a task that is intrinsically NP-hard, is a fundamental challenge in the fields of optimization, machine learning, image processing, and control theory. One popular method is the relaxed approximation method, for example, convex relaxed method [20, 21, 22], nonconvex relaxed method [23, 24] and DC (difference of convex functions) relaxed method [25, 26]. Due to an enhanced comprehension of the variational characteristics associated with the $\ell_{0}$-norm, numerous studies have addressed sparse optimization by directly optimizing the $\ell_{0}$-norm [27, 28]. Additionally, successive coordinatewise convex problem can approximate $\ell_{0}$-norm penalty problem [29].
@@ 1.2 Contributions
1) The optimal sparse feedback gain of the considered LQ problem can be represented by the global minimizer of a constrained optimization problem with sparsity penalty. Compared with [40, 17, 14], the feedback gain with general sparsity structure is investigated rather than the block-diagonal feedback gain, and instead of the infinite-dimensional optimization that is solved by finite-dimensional approximation method [41, 42, 43], the proposed optimization problems of this paper is finite-dimensional. Furthermore, in contrast to [15, 4, 44, 16], the algorithms introduced in this paper provide an optimality and stabilizing guarantee rather than heuristic algorithm or stationary-point-guarantee.
2) The results of this paper is of independent interest in optimization field. Note that the core algorithm used to solve the constrained convex optimization problem with weighted-$\ell_{1}$ sparsity promoting penalty is the one of [39], and whereas the methodology of [39] cannot be used directly due to the presence of a complicated subproblem, necessitating the development of an iterative algorithm based on BSUM methodology. To highlight that the BSUM framework consists of successive solving of subproblems induced by auxiliary functions, we design special auxiliary functions such that each subproblem exhibits a closed-form solution. Furthermore, we have established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function. To our best knowledge, this phenomenon has not been reported in the field of sparse optimization.
3) The variational properties of $\ell_{0}$-penalty is firstly studied for decentralized LQ problem. Motivated by the convex relaxation framework of [29], we show that the $\ell_{1}$-penalty induced problems can be seen as approximations of $\ell_{0}$-penalty induced problem in the field of decentralized LQ optimal control. Noting that [29] studies an unconstrained smooth optimization problem, the considered problem of this paper is constrained and nonsmooth. Hence, the conclusion and techniques of proof presented in our analysis do not exhibit parallel generalizations of those of [29]. Moreover, the interconnection between the proposed optimization problems is illustrated in Fig. 1 below.
The rest of this paper is organized as follows. In Section 2, the basic formulations and assumptions of decentralized LQ problem are proposed. In Section 3, we will introduce a two-timescale algorithm framework based on primal-dual splitting method to solve the weighted-$\ell_{1}$-sparsity promoting optimization problem, which encompasses a complicated subproblem. Such a subproblem is investigated in Section 4. Section 5 provides the convergence analysis of the proposed algorithm and studies quadratic sparsity promoting optimization problem, which can be regarded as an acceleration of weighted-$\ell_{1}$ promoting problem. In Section 6, we will study an $\ell_{0}$-sparsity promoting optimization problem which is nonconvex and discontinuous, and propose some theorems to characterize its variational properties.
$\ell_{0}$-sparsity promoting Section 6 weighted-$\ell_{1}$-sparsity promoting Section 3, 4, 5 quadratic sparsity promoting Section 5 coordinate convex problem Section 6 subproblem 1 $\Leftrightarrow$ weighted-$\ell_{1}$-sparsity promoting, subproblem 2 $\Leftrightarrow$ closed-form solution convex relaxationconvex relaxationaccelerationapproximationBSUM method
@@
$\mathbf{prox}_{\tau g}^{\mathcal{Y}}(z):=\underset{y\in\mathcal{Y}}{\operatorname*{argmin}}\left\{g(y)+\frac{1}{2\tau}\left\|y-z\right\|^{2}\right\},\quad\forall z\in\mathbb{R}^{n}.$
For any $\alpha\in\mathbb{R}$, the $\alpha$-level set of a function $f:\mathbb{R}^{n}\to[-\infty,\infty]$ is the set ${\rm lev}_{\leq\alpha}f=\{x\in\mathbb{R}^{n}\colon f(x)\leq\alpha\}.$ We introduce the notation $M\lesssim N$, meaning $M\leq CN$ with some generic bounded constant $C>0$. For any subset $C\subseteq\mathbb{R}^{n}$, the indicator function of $C$ is defined to be the extended real-valued function given by
$\delta_{C}(x)=\left\{\begin{aligned} 0,~{}~{}&x\in C,\\ \infty,~{}~{}&x\notin C.\end{aligned}\right.$
@@ 2 Formulation
We introduce the following notations with $p=m+n$
$F=\begin{bmatrix}A&B_{2}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p},~{}G=\begin{bmatrix}0\\ I_{m}\end{bmatrix}\in\mathbb{R}^{p\times m},$
$Q=\begin{bmatrix}B_{1}B_{1}^{\top}&0\\ 0&0\end{bmatrix}\in\mathbb{S}^{p},~{}R=\begin{bmatrix}C^{\top}C&0\\ 0&D^{\top}D\end{bmatrix}\in\mathbb{S}^{p}.$
@@
$W=\begin{bmatrix}W_{1}&W_{2}\\ W_{2}^{\top}&W_{3}\end{bmatrix}\in\mathbb{S}^{p}$
with $W_{1}\in\mathbb{S}_{++}^{n},W_{2}\in\mathbb{R}^{n\times m},W_{3}\in\mathbb{S}^{m}$. Regard $\Theta_{i}(W)=F_{i}W+WF_{i}^{\top}+Q$ as a block matrix, i.e.,
@@
The following theorem introduces a subset of stabilizing controller gains, and finds an upper bound of performance criterion (2).
@@
and let $\mathscr{K}=\{K=W_{2}^{\top}W_{1}^{-1}\colon W\in\mathscr{C}\}$. Then,
@@
    $K\in\mathscr{K}$ stabilizes the closed-loop system;
@@
    $\langle R,W\rangle\geq\|H_{i}(s)\|_{\mathcal{H}_{2}}^{2},~{}~{}i=1,\dots,M,$
@@
Proof: The proof is familiar with that of Theorem 1 of [40], but for the sake of completeness, we will include short proof here. For statement 1, $\Theta_{i,1}\preceq 0$ implies that
$(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})W_{1}$ $+W_{1}(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})^{\top}$ (3)
$+B_{1}B_{1}^{\top}\preceq 0.$
@@
$\langle R,W\rangle$ $={\rm Tr}(CW_{1}C^{\top}+DW_{3}D^{\top})$
$\geq{\rm Tr}(CW_{c,i}C^{\top}+DKW_{c,i}K^{\top}D^{\top})$
$={\rm Tr}((C-DK)W_{c,i}(C-DK)^{\top}).$
[…]
@@
If $z(t)$ exhibits the following form which is studied in [14]
$z(t)=\begin{bmatrix}Q^{\frac{1}{2}}\\ 0\end{bmatrix}x(t)+\begin{bmatrix}0\\ R^{\frac{1}{2}}\end{bmatrix}u(t)$
with $Q,R\succ 0$, then $C^{\top}C\succ 0$, $D^{\top}D\succ 0$ and $C^{\top}D=0$ naturally hold. In fact, when $F$ is known, the classic infinite-horizon LQ problem
$\min_{K\in\mathbb{R}^{m\times n}}{}{}$ $J(K)$ (4)
${\rm s.t.}{}{}$ $K\in\mathcal{S}$
with stabilizing feedback gain set $\mathcal{S}$ is equivalent to the following convex optimization problem [48]
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle$ (5)
${\rm s.t.}{}{}$ $W\in\mathscr{C}.$
However, we want to expand the range of applicability in cases where $F$ is unknown but convex-bounded; see Assumption 2. In this case, by Theorem 1, the optimization problem (5) is no longer equivalent to (4), but the objective function of (5) is the upper bound of $J(K)$.
In this study, we focus on a general decentralized control problem, assuming that the feedback gain exhibits a general sparsity structure rather than a block diagonal structure [40]. The fundamental idea in our paper is that the sparse stabilizing feedback gain, regardless of its structure, can be achieved by the design of a diagonal matrix $W_{1}$ and a sparse matrix $W_{2}$ with some sparse structure. Then, feedback gain $K=W_{2}^{\top}W_{1}^{-1}$ can preserve the sparsity structure of $W_{2}$ and stabilize the closed-loop system.
@@
In this paper, we will use the weighted-$\ell_{1}$-norm, the piecewise quadratic function, and the $\ell_{0}$-norm as sparsity promoting functions, which correspond to different optimization problems. Letting $N=n(n-1)/2$, $e_{i}=(\mathbf{0}_{i-1}^{\top},1,\mathbf{0}_{p-i}^{\top})^{\top}\in\mathbb{R}^{p}$,
$E_{i}^{j}=\begin{bmatrix}e_{i}^{\top}\\ \vdots\\ e_{i}^{\top}\end{bmatrix}\in\mathbb{R}^{j\times p},~{}~{}\hat{E}_{i}^{j}=\begin{bmatrix}e_{i}\\ \vdots\\ e_{j}\end{bmatrix}\in\mathbb{R}^{p(j-i+1)},$
$V_{*1}=\begin{bmatrix}V_{11}\\ \vdots\\ V_{N1}\end{bmatrix}\in\mathbb{R}^{N\times p},~{}~{}V_{*2}=\begin{bmatrix}V_{12}\\ \vdots\\ V_{N2}\end{bmatrix}\in\mathbb{R}^{Np},$
we denote
$V_{*1}=\begin{bmatrix}E_{1}^{n-1}\\ E_{2}^{n-2}\\ \vdots\\ E_{n-1}^{1}\end{bmatrix},~{}~{}V_{*2}=\begin{bmatrix}\hat{E}_{2}^{n}\\ \hat{E}_{3}^{n}\\ \vdots\\ \hat{E}_{n}^{n}\end{bmatrix}.$ (6)
Then, for the weighted-$\ell_{1}$-norm of a matrix $P$
$g(P)=\sum_{i,j}w_{ij}|P_{ij}|$ (7)
with $w_{ij}>0$ for all $i,j$, we introduce the following optimization problem
$\min_{W,P}~{}~{}$ $\langle R,W\rangle+\gamma g(P)$ (8)
${\rm s.t.}~{}~{}$ $W\in\mathbb{S}^{p}_{+},$ (9)
$\Psi_{i}\in\mathbb{S}_{+}^{n},\forall i=1,\dots,M,$ (10)
$-V_{j1}WV_{j2}=0,\forall j=1,\dots,N,$ (11)
$V_{1}WV_{2}^{\top}-P=0$ (12)
with $\gamma>0$, $V_{1}=[0,I_{m}],V_{2}=[I_{n},0]$ and
@@
Remark 2.
The above problem is the core optimization problem of this paper, where the constraints (9) and (10) are equivalent to $W\in\mathscr{C}$. Additionally, by the definitions of $V_{*1}$ and $V_{*2}$, the constraint (11) implies that $W_{1}$ is a diagonal matrix, and the constraint (12) is equivalent to $P=W^{\top}_{2}$ so that sparsity penalty term $\gamma g(P)=\gamma g(W^{\top}_{2})$. Together with constraints (9)-(12) and Theorem 1, the matrix $K=W_{2}^{\top}W_{1}^{-1}$ is a stabilizing feedback gain with general sparsity structure.
Remark 3.
In fact, let
$D_{k}=\begin{bmatrix}0\\ I_{k}\end{bmatrix}\in\mathbb{R}^{n\times k},$
then the $N$ equality constraints
$V_{j1}WV_{j2}=0,~{}j=1,\dots,N$ (13)
can be converted into $n-1$ equality constraint:
$e_{k}V_{2}WV_{2}^{\top}D_{n-k}=0,~{}k=1,\dots,n-1$
with $e_{k}=(0,\dots,1,\dots,0)$. For the convenience and simplicity of formulas’ deducing, we still use $N$ equality constraints (i.e., (11)) to describe the diagonal property of matrix $W_{1}$.
Remark 4.
The group-sparsity promoting function can be defined likewise. For example, letting the matrix $P\in\mathbb{R}^{m\times n}$ be partitioned into submatrices $P_{ij}\in\mathbb{R}^{m_{i}\times n_{j}}$, the group weighted-$\ell_{1}$-norm can be defined as
$g_{\rm group}(P)=\sum_{i,j}w_{ij}\|P_{ij}\|_{F}.$
By replacing $\gamma g(P)$ of (8) by $\gamma g_{\rm group}(P)$ and modifying the constraint (11) such that $W_{1}$ is a block-diagonal matrix with proper size, we then obtain a group sparse feedback gains $K=W_{2}^{\top}W_{1}^{-1}$. Remarkably, all the results of this paper could be generalized into the group-sparsity-promoting setting, and for the sake of simplicity we only focus on how to obtain a sparse (instead of group sparse) feedback gain in the following sections.
Remark 5.
Generally, a sparse (or group-sparse) feedback gain $K$ can lead to a multi-agent decentralized control system. For example, given a group-sparse feedback gain
$K=\begin{bmatrix}K_{11}&\cdots&K_{1q}\\ \vdots&\ddots&\vdots\\ K_{p1}&\cdots&K_{pq}\end{bmatrix}\in\mathbb{R}^{m\times n}$
with $K_{ij}\in\mathbb{R}^{m_{i}\times n_{j}}$ and $\sum_{i=1}^{p}m_{i}=m,\sum_{j=1}^{q}n_{j}=n$, it can induce a $p$-agents decentralized system. Specifically, the global control action is composed of local control actions: $u(t)=[u_{1}(t)^{\top},\dots,u_{p}(t)^{\top}]^{\top}$, where $u_{i}(t)\in\mathbb{R}^{m_{i}}$ is the control input of agent $i$. At time $t$, agent $i$ directly observes a partial state $x_{\mathcal{I}_{i}}(t)$, and $\mathcal{I}_{i}$ is given by
$\mathcal{I}_{i}=\bigcup_{j=1,\dots,q,~{}K_{ij}\neq 0}\left\{n_{j-1}+1,\dots,n_{j}\right\}$
with $n_{0}=0$. Here, $\mathcal{I}_{i}$ is a fixed subset of $[n]\doteq\{1,\dots,n\}$ and $x_{\mathcal{I}_{i}}(t)$ denotes the subvector of $x(t)$ with indices in $\mathcal{I}_{i}$. That is why in this paper we investigate the decentralized LQ problem by studying optimization problem
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle+\gamma s(V_{1}WV_{2}^{\top})$ (14)
${\rm s.t.}{}{}$ $W\in\mathscr{C}$
with sparsity promoting function $s(\cdot)$. It is important to note that the $p$-agents decentralized systems in [40, 17, 14] satisfy
$\mathcal{I}_{i}\cap\mathcal{I}_{j}=\emptyset,~{}~{}\forall i,j\in[p],i\neq j.$
However, $\mathcal{I}_{i}$ and $\mathcal{I}_{j}$ may overlap in our setting, and hence the considered problem of this paper is a generalization of those of [40, 17, 14].
Optimization problem (8)-(12) can be equivalently expressed in a compact form
$\min_{W,P}{}{}$ $\langle R,W\rangle+\gamma g(P)$ (15)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
@@
and convex cone $\mathcal{K}$ is defined as $\mathcal{K}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times 0\times\cdots\times 0.$ Since the positive semi-definite cone is self-dual, it is straightforward to express the dual cone $\mathcal{K}^{*}$ as $\mathcal{K}^{*}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times\mathbb{R}\times\cdots\times\mathbb{R}.$